### v1

# Labels or Input? Rethinking Augmentation in Multimodal Hate Detection

Abstract
The modern web is saturated with multimodal content, intensifying the challenge of detecting hateful memes, where harmful intent is often conveyed through subtle interactions between text and image under the guise of humor or satire. While recent advances in Vision-Language Models (VLMs) show promise, these models lack support for fine-grained supervision and remain susceptible to implicit hate speech. In this paper, we present a dual-pronged approach to improve multimodal hate detection. First, we propose a prompt optimization framework that systematically varies prompt structure, supervision granularity, and training modality. We show that prompt design and label scaling both influence performance, with structured prompts improving robustness even in small models, and InternVL2 achieving the best F1-scores across binary and scaled settings. Second, we introduce a multimodal data augmentation pipeline that generates 2,479 counterfactually neutral memes by isolating and rewriting the hateful modality. This pipeline, powered by a multi-agent LLM–VLM setup, successfully reduces spurious correlations and improves classifier generalization. Our approaches inspire new directions for building synthetic data to train robust and fair vision-language models. Our findings demonstrate that prompt structure and data composition are as critical as model size, and that targeted augmentation can support more trustworthy and context-sensitive hate detection.
Codehttps://github.com/sahajps/Labels-or-Input
Datasetshttps://hf.co/datasets/sahajps/Meme-Sanity
Introduction
The rise of Vision-Language Models (VLMs) such as CLIP (Radford et al. 2021), Flamingo (Alayrac et al. 2022), and InternVL2 (Chen et al. 2024) has led to significant progress in understanding and processing multimodal content. These models are increasingly applied to downstream tasks, including meme classification and online content moderation. At the same time, visual harmful content has surged across platforms like X and Instagram, where images and videos constitute between 30% of shared content on text-centric platforms and nearly 100% on image-first platforms (Peng, Lu, and Shen 2023; Yang, Davis, and Hindman 2023; Pfeffer et al. 2023). This shift underscores the growing need for scalable, robust methods that can effectively address the complexities of multimodal hate speech detection in contemporary social media environments (González-Aguilar, Segado-Boj, and Makhortykh 2023; Heley, Gaysynsky, and King 2022; Solea and Sugiura 2023). Moreover, recent studies find that VLMs are more susceptible to harmful content than their unimodal counterparts, with image-text combinations inducing unsafe behaviors even under benign conditions (Hee and Lee 2025).
While prior work has addressed aspects of hate meme detection, many existing approaches focus on fine-tuning pre-trained models without exploring how prompting might be used to enrich supervision or guide classification (Lippe et al. 2020; Hermida and Santos 2023). Although prompt engineering has shown promise in unimodal tasks as a lightweight method to influence model behavior (Furniturewala et al. 2024), its utility in the multimodal setting, particularly for enhancing training labels or shaping classification decisions in VLMs, remains underexplored. These research gaps motivate our core research question: Is improving prompting or refining the underlying data more effective than a standard baseline in multimodal hateful content classification?
Figure 1: Overview of the augmentation pipeline. Red and green boxes indicate hateful and non-hateful content. In part B, the augmented Data is created for the green-red pairs – cases where the background is non-hateful but is used in combination with a hateful caption, resulting in a hateful interpretation.
Recent studies also highlight significant limitations in dataset quality and data augmentation practices. Several datasets are constructed through manual pipelines that involve overlaying text on stock images, yielding templated meme formats (Nandi, Sharma, and Chakraborty 2024). Broader surveys on multimodal hate content (Hee et al. 2024) emphasize the growing presence of hateful speech across text, image, and video, but often stop short of proposing actionable modeling strategies. Synthetic augmentation methods such as MixGen (Hao et al. 2023) attempt to increase diversity by randomly mixing images and captions, but frequently produce semantically incoherent pairings that reduce realism. In parallel, cultural and representational biases, and stereotypes embedded in VLMs (Madasu, Lal, and Howard 2025; Liu et al. 2025) are increasingly recognized yet get very little attention in multimodal hate speech detection, especially culturally grounded hateful memes (Bui, von der Wense, and Lauscher 2025).
In this study, we investigate two complementary strategies (see Figure 1) for improving hateful meme classification: (1) prompt optimization, which leverages structured prompting to enrich supervision across modalities and label formats; and (2) multimodal augmentation, which generates counterfactually neutral memes to reduce spurious correlations.
We conduct experiments on the Facebook Hateful Memes dataset (Kiela et al. 2020), which contains over 10K image-caption pairs labeled for binary hatefulness. Our evaluation involves fine-tuning both unimodal and multimodal models under each strategy. This setup enables a direct comparison of prompt-informed supervision and data-centric augmentation with standard baselines for robust multimodal classification. Our study finds that label augmentation, such as scaling hatefulness, shows consistent patterns across modalities and prompting styles. This indicates that well-augmented multimodal hate labels can enable a more nuanced understanding of harmful content beyond simple binary classification. We also observe that input-level multimodal augmentation improves performance across modalities when fine-tuning pretrained models. Human evaluation further supports the quality of the augmented data, highlighting its potential in reducing toxicity and pointing to the broader promise of large multimodal models when used strategically in multi-agent systems.
Contributions.
In line with our ethical commitment, we deliberately avoid generating new hateful content in the name of multimodal data augmentation. The prevalence of such material on social media is already deeply concerning, and our goal is to mitigate harm, not amplify it. In short, this work makes the following key contributions:
  • We extend the Facebook Hateful Memes dataset by generating 2479 counterfactually neutral memes, offering a novel benchmark for evaluating robustness to spurious text-image correlations.
  • We introduce a structured prompt optimization framework for multimodal classification, demonstrating that prompt design—long studied in language-only settings—has overlooked yet significant effects in vision-language tasks.
  • We explore multimodal augmentation as a debiasing strategy, providing a proof-of-concept pipeline that could generalize to related domains such as cultural bias mitigation in VLMs.
Related Work
Humor and satire are frequently used to mask derogatory content or reinforce stereotypes while maintaining plausible deniability (Smuts 2010). As noted in prior work, this form of disparagement humor can reduce social sanctions for expressing prejudice and foster political intolerance (Perks 2012; Ferguson and Ford 2008; Masood et al. 2024), provoking reactive sharing behaviors that reinforce in-group superiority and out-group hostility (Masood et al. 2024). These dynamics make hatefulness in memes both culturally contingent and pragmatically difficult to detect through surface features alone.
To address the difficulty of interpreting memes, prior work has shifted from unimodal to multimodal approaches. Early efforts relied on isolated text or image classifiers, but these often failed to capture the semantic interplay that drives hateful meaning in memes (Lippe et al. 2020). The release of the Hateful Memes dataset (Kiela et al. 2020) catalyzed the development of multimodal fusion models capable of integrating visual and linguistic features.
Subsequent models incorporated pre-trained vision-language architectures such as CLIP (Radford et al. 2021) and Flamingo (Alayrac et al. 2022), enabling improved joint reasoning over text and image inputs. However, many of these systems still struggle with out-of-distribution examples and culturally specific memes, which are underrepresented in benchmark datasets (Bui, von der Wense, and Lauscher 2025; Madasu, Lal, and Howard 2025). Vision-language models exhibit known cultural and representational biases, which can be especially harmful in hate detection contexts. These biases stem from skewed pretraining distributions and insufficient coverage of minority identities or cultural idioms (Liu et al. 2025; Madasu, Lal, and Howard 2025). Additionally, VLMs are more prone than unimodal models to unsafe behaviors triggered by multimodal inputs (Liu et al. 2024).
Such vulnerabilities highlight the need for both architectural and data-centric solutions. Our contributions involving prompt optimization and semantically coherent augmentation function as lightweight forms of regularization that improve interpretability and mitigate failure modes related to spurious correlations, stereotype reinforcement, and brittle visual-textual associations.
In summary, the need for improved multimodal hate speech detection has motivated two complementary research directions: using prompting to better structure supervision and labels, and using input data augmentation to improve robustness.
Prompting as a mechanism for supervision.
While prompting has become an effective tool for adapting large language models (Furniturewala et al. 2024), its applications in multimodal settings are less explored (Wu et al. 2024). Existing VLMs rarely leverage structured prompts to guide classification decisions or clarify task framing. Yet, the presence or absence of hateful intent often hinges on subtle shifts in phrasing or cultural markers. Our approach introduces task- and label-specific prompt templates to serve as soft supervision, helping models differentiate between hateful and non-hateful combinations of text and image, particularly in ambiguous or sarcastic cases.
Augmentation as a tool for improving robustness.
In parallel, prior work highlights the limitations of hate meme datasets built from templated or synthetic formats (Nandi, Sharma, and Chakraborty 2024), and the failure of augmentation techniques like MixGen (Hao et al. 2023) to preserve semantic coherence. On the other hand, some studies (Koushik et al. 2025; Gu et al. 2025) use the textual direction of data augmentation by adding extra context (such as a detailed description of the image, CoT reasoning process, etc.). While converting inputs to a single modality (such as text) can improve hate tagging, this approach is not particularly effective for debiasing memes during fine-tuning. The training process may still cause the model to associate certain labels with specific image component words like “Muslim” or “Black”. To address this, we adopt a counterfactual data augmentation strategy that constructs minimally modified meme variants, e.g., replacing hateful captions with neutral ones while retaining visual features. This helps models disentangle surface form from semantic content, reducing overfitting to visual cues and improving generalization to culturally grounded memes.
Methodology
We evaluate two distinct strategies for improving hate meme classification: (1) prompt optimization, which treats prompting as a multivariate design space across learning modality, structure, label format, and fine-tuning; and (2) multimodal augmentation, which seeks to improve model robustness by systematically neutralizing hateful captions while preserving visual context. These approaches are compared to assess their effectiveness in enhancing classification accuracy, robustness, and fairness.
Prompt Optimization Framework
For our first set of experiments, we explore a full-factorial design spanning 3 learning styles (text-only: fine-tuning, multimodal: fine-tuning & prompting), 2 prompt strategies (simple vs. category-based), and 2 output types (binary vs. scale-based). We use prompting as a labeling instrument to generate enriched supervision signals. Specifically, we design structured prompt templates - ranging from simple to category-specific prompts - and from binary to scale-based outputs, and use them with a teacher model (GPT-4o-mini [^1] https://platform.openai.com/docs/models/gpt-4o-mini) on a subset of training data (due to high cost associated with the use of GPT-4o-mini with vision inputs) to support scale-based labeling and generate interval annotations (0–9), thereby enhancing label granularity. These enhanced labels are then used to fine-tune both a large multimodal model (InternVL2 8B (Chen et al. 2024)) and a smaller unimodal model (DistilBERT (Sanh et al. 2020)), allowing us to evaluate the impact of prompt-informed supervision. All prompt variations are carefully controlled and documented to ensure consistency (refer to supplementary information (SI)).
Problem formulation
Let each input meme be defined as a tuple $(I_{i},T_{i},y_{i})$, where $I_{i}$ is the image, $T_{i}$ is the caption text, and $y_{i}\in\{0,1\}$ is the binary ground-truth hatefulness label. Given a model $M$ parameterized by $\theta$, a prompt template $P$, and a label format $L$, the model’s predicted label $\hat{y}_{i}$ is generated as:
$\hat{y}_{i}=M(I_{i},T_{i};\theta,P,L)$
Prompt Construction
We define two strategies:
  • Simple Prompt ($P_{s}$): A direct instruction that asks whether the meme is hateful, without further elaboration.
  • Category Prompt ($P_{c}$): A structured prompt that defines subtypes of hate (e.g., misogyny, xenophobia, political hate), encouraging the model to attend to specific cues.
All prompt templates are modular and composable with both label formats (see details in SI). We fix prompt wording and structure across conditions to minimize confounds.
Label Format Design
Two types of outputs are evaluated:
  • Binary Output ($L_{b}$): The model is learning to tag hatefulness label as TRUE or FALSE.
  • Scale Output ($L_{s}$): The model is getting trained on a hatefulness score on an integer scale from 0 (not hateful) to 9 (extremely hateful). And later, at the test time, it is being mapped to TRUE (5-9) or FALSE (0-4).
To enable fine-tuning on scale outputs, we use a teacher model to generate soft labels for training. Incorrect labels (i.e., those inconsistent with the binary ground truth) are manually filtered out to ensure label quality.
Ultimately, this results in a $3\times 2\times 2$ factorial experiment across learning modality ($M$), prompt structure ($P$), and label format ($L$) with 12 distinct configurations. Each configuration is evaluated independently to isolate the marginal contributions of each component and to test for interaction effects between prompting and fine-tuning.
Multimodal Augmentation
For our second set of experiments, we explore a data-centric augmentation approach, applying a multi-step and multi-agent pipeline to regenerate non-hateful variants of memes using LLM and VLM tools. This approach produces a bias-aware augmented dataset designed to reduce spurious correlations while preserving semantic and visual coherence.
Problem Formulation
Let $D=\{(I_{i},T_{i},y_{i})\}_{i=1}^{N}$ denote a dataset of memes, where $I_{i}$ is the image, $T_{i}=[w_{1},\ldots,w_{|T_{i}|}]$ is the tokenized caption text, and $y_{i}\in\{0,1\}$ is the binary hatefulness label. We define a counterfactual data augmentation objective to construct neutral variants $\tilde{M}_{i}=(\tilde{I}_{i},\tilde{T}_{i},\tilde{y}_{i})$ by replacing hateful captions $T_{i}$ with context-preserving neutral captions $\tilde{T}_{i}$, where $\tilde{y}_{i}=0$.
This augmentation targets memes for which the caption is the primary hateful modality—i.e., where $I_{i}$ is semantically benign. The hypothesis is that injecting such counterfactuals will decorrelate textual toxicity from visual features and improve generalization.
Hatefulness Attribution
To identify the modality responsible for hatefulness, we define a modality-level attribution function:
$H_{i}=f_{\phi}(g_{\psi}(I_{i}),T_{i})\in\{\texttt{image},\texttt{text},\texttt{both},\texttt{none}\}$
Where $f_{\phi}$ is a zero-shot classifier instantiated via Qwen2.5-14B (Yang et al. 2025) and $g_{\psi}$ is defined in the forthcoming paragraph. Only samples with $H_{i}=\texttt{text}$ are selected for augmentation. This procedure ensures that alterations are made only to memes where the caption is the dominant hateful component.
Context-Preserving Caption Neutralization
We first generate a background description $D_{i}$ for the image $I_{i}$ using a vision-language model $g_{\psi}$ (InternLM-XComposer-2.5-7B (Zhang et al. 2024)): $D_{i}=g_{\psi}(I_{i})$
The model is constrained to exclude any recognition of overlaid text. Given the original caption $T_{i}$ and image context $D_{i}$, we generate a rewritten, neutral caption $\tilde{T}_{i}$ using a generative model $h_{\omega}$ (GPT-4o-mini):
$\tilde{T}_{i}=h_{\omega}(T_{i},D_{i})\text{, such that }\texttt{Hate}(\tilde{T}_{i})=0$
We enforce semantic coherence and topical grounding through conditioning on $D_{i}$. While $h_{\omega}$ includes built-in safety mechanisms to promote safe responses, we still assess its safety performance as part of our human evaluation on a subset of the augmented multimodal dataset.
Meme Regeneration and Dataset Augmentation
We synthesize a regenerated meme $\tilde{M}_{i}=(\tilde{I}_{i},\tilde{T}_{i},0)$ using a rendering function $r_{\gamma}$ implemented via Gemini 2.0 Flash (experimental) [^2] https://cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/2-0-flash, where, $\tilde{I}_{i}=r_{\gamma}(D_{i},\tilde{T}_{i})$. This model overlays $\tilde{T}_{i}$ on $I_{i}$’s equivalent using layout-preserving visual information. The final augmented dataset is defined as:
$\tilde{D}=\{\tilde{M}_{i}=(\tilde{I}_{i},\tilde{T}_{i},0)\mid H_{i}=\texttt{text},\texttt{Hate}(\tilde{T}_{i})=0\}$
Where $g_{\psi}(\tilde{I}_{i})\approx g_{\psi}(I_{i})$, i.e., the background descriptions of old and newly augmented memes are similar.
This corpus $\tilde{D}$ is merged with the original dataset $D$ for downstream training. We evaluate its impact on accuracy, fairness, and out-of-distribution robustness. Table 1 shows some visual examples for the pipeline.
Meme Typology and Target Class
We define four meme types based on modality-specific hatefulness: HH (hateful image & hateful caption), HN (hateful image & non-hateful caption), NH (non-hateful image & hateful caption), and NN (non-hateful image & caption).
Empirical analysis indicates that the majority of hateful memes belong to the NH category. Our augmentation strategy explicitly targets this class to reduce spurious associations between benign visual features and toxic labels.
Full implementation details, including model prompts and architecture-specific settings, are provided in the SI.
Empirical Evaluation
We now describe the datasets, baselines, and models used to evaluate our two experimental approaches, prompt optimization and multimodal augmentation, for the task of classifying hate memes. Our evaluation focuses on classification accuracy and weighted-F1 score, as standard in the Facebook Hateful Memes benchmark. Note that in each setting, we are splitting the training set into two parts: one for training and one for validation. For testing, we are combining both seen and unseen test data. Additional details, including prompt templates and training configurations, are provided in the SI.
Example 1 Example 2 Example 3 Example 4
Hateful Meme
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Non-hateful Augmentation
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Example 5 Example 6 Example 7 Example 8
Hateful Meme
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Non-hateful Augmentation
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 1: Input-output examples from multimodal data augmentation pipeline.
Datasets
We use the Facebook Hateful Memes dataset (Kiela et al. 2020), a benchmark corpus for multimodal hate speech detection. The dataset consists of over 10,000 image-caption pairs for training and 3,000 for test evaluation. Each meme is annotated with a binary label indicating whether it is hateful. Captions are written in natural language and paired with images that range from benign stock photos to visually suggestive content.
For the prompt optimization experiments, we use the original dataset labels (binary) and augment them with scale-based annotations using GPT-4o-mini as a teacher model.
For the multimodal augmentation experiments, we extract a subset of the training set in which hatefulness is localized to the caption only (i.e., image visuals are neutral). These samples are regenerated using our multi-step and multi-agent pipeline to create counterfactually non-hateful memes while preserving visual context.
Fine-Tuning Models and Setup
To evaluate both label augmentation and multimodal data augmentation strategies, we fine-tune (sometimes prompt, where applicable, such as InternVL2) a suite of unimodal and multimodal models on the Facebook Hateful Memes dataset. For fine-tuning, AdamW is used as an optimizer with cross-entropy as a loss function, which is standard in classification tasks.
(a) Prompt Optimization Experiments
For the prompt optimization track, we evaluate 12 experimental configurations from a $3\times 2\times 2$ factorial design varying:
  • Learning Type: using text (fine-tuning), multimodal (fine-tuning and prompting);
  • Prompt Style: simple vs. category-based;
  • Label Format: binary vs. scale-based labels for learning.
We compare performance across zero-shot prompting with InternVL2 and prompt-informed fine-tuning, using InternVL2 and DistilBERT. All smaller models are fully fine-tuned with the same hyperparameters to ensure comparability. For InternVL2, we apply LoRA (Hu et al. 2022) -based parameter-efficient tuning. The learning rate and batch size are taken as $10^{-4}$ and $16$ respectively. Results and model sizes are given in Table 2.
(b) Multimodal Augmentation Experiments
For this augmentation track, we fine-tune each pre-trained model on both the original dataset and an extended version $\tilde{D}$, which includes counterfactually neutral memes. We experiment with three model types: unimodal (text), unimodal (vision), and multimodal (text + vision), allowing a direct comparison of robustness to spurious correlations. To address class imbalance, we apply downsampling. Fine-tuning is performed on the last two layers of the encoders, with a learning rate of $2\times 10^{-5}$ and batch size of 128. Table 3 presents the accuracy and F1-scores for all models.
Models used are publicly available checkpoints or APIs. Hyperparameters are also kept consistent across settings to control for confounding effects. Full training and inference configurations are included in the SI.
Category Model Size Prompt Training dataset
Original (Binary) Scaled
Accuracy F1-score Accuracy F1-score
Unimodal
(fine-tuning)
[…]
Multimodal
[…]
Multimodal
(fine-tuning)
[…]
Table 2: Classification results with binary and scaled labels on simple and category styled prompts. The original dataset has the binary labels. Scaled labels were granular values generated by a teacher model, GPT-4o-mini.
Category Model Size Learning Training dataset
Original (Binary) Extended
Accuracy F1-score Accuracy F1-score
Unimodal
(fine-tuning)
[…]
Unimodal
(fine-tuning)
[…]
Multimodal
(fine-tuning)
[…]
Table 3: Classification results with and without newly augmented non-hateful meme data. Note: Until specified, we have used the base variant of the models mentioned. Note that the size is written as an approximate number of parameters in the model.
Results and Analysis
We present the results and implications of our data augmentation strategy using binary and scaled labels, evaluated via prompting and fine-tuning. While we report both accuracy and weighted F1-score, we focus on F1-score due to its greater sensitivity in class-imbalanced classification tasks.
Effectiveness of Prompt Optimization
We evaluate two prompting strategies—simple and category—across unimodal and multimodal models, under both binary and scaled label regimes (Table 2).
Prompting: Category vs. Simple
Under the binary label setting, the category prompt improves performance for multimodal prompting with InternVL2 (F1: 63.00 vs. 61.36), and further boosts it under fine-tuning (F1: 68.30 vs. 67.51). For scaled labels, results are mixed: a slight decrease with prompting and a minor increase (F1: 64.74 vs. 64.63) with fine-tuning. Since the scaled dataset is a small subset of the full training data, these improvements may grow with larger datasets.
Surprisingly, for DistilBERT, the category prompt underperforms the simple prompt in both cases (F1: 49.46 vs. 59.87 and 55.58 vs. 58.41), likely due to overloading the smaller model with excessive category information.
Fine-tuning vs. Prompting
Multimodal fine-tuning with InternVL2 achieves the highest overall F1-scores: 68.30 (binary) and 64.74 (scaled), outperforming prompting-only setups across the board. This highlights the advantage of full model adaptation for downstream hatefulness classification. Even so, zero-shot prompting with InternVL2 yields competitive performance (F1: 63.00 binary; 63.89 scaled), indicating strong generalization capabilities with appropriate prompts.
Implications
These trends reveal that:
  • Prompt quality matters: Structured, informative prompts (category) improve performance in most cases, particularly for larger models.
  • Scaled labels are learnable: Despite added complexity, models trained on scaled annotations achieve robust performance (e.g., 64.74 F1 for InternVL2), enabling nuanced hatefulness detection.
  • Lightweight deployment is viable: Even DistilBERT achieves 55.58 F1 on scaled labels with category prompts, showing promise for low-resource or real-time applications.
Effectiveness of Multimodal Data Augmentation
To assess the value of augmenting our training set with counterfactually neutral memes, we fine-tune a diverse range of popular unimodal and multimodal models, including BERT (Devlin et al. 2019), RoBERTa (Liu et al. 2019), ALBERT (Lan et al. 2020), ViT (Wu et al. 2020), BEiT (Bao et al. 2022), DeiT (Touvron et al. 2021), and CLIP (Radford et al. 2021), using identical hyperparameters across both the original and extended datasets. We aim to determine whether exposure to visually and lexically similar non-hateful memes improves model generalization and robustness.
F1 Gains Across Modalities
Across most model types, the extended dataset yields modest gains in accuracy (typically <<1%), but more notable improvements in F1-score, indicating better handling of class imbalance and edge cases. For example:
  • BERT improves from 55.50 to 57.95 in F1 (+2.45), with a negligible drop in accuracy (59.08 to 58.20).
  • RoBERTa shows a larger relative F1 gain, from 47.18 to 52.54 (+5.36), despite a slight accuracy decrease.
  • CLIP, the sole multimodal model, is evaluated, improving from 47.71 to 51.77 in F1 (+4.06), showing that the augmentation is also effective in the joint vision-language space.
The only unimodal model that sees near-constant performance is ALBERT, which improves slightly in F1 from 54.96 to 58.54 (+3.58), but with no meaningful change in accuracy. Among vision-only models, BEiT shows the best improvement (F1: 53.38 to 55.43), while DeiT lags with marginal change.
Text Dominance and Modality Gaps
As expected, text-based models outperform vision-only models across both datasets. For instance, on the extended data, BERT and RoBERTa achieve higher F1-scores (57.95 and 52.54, respectively) compared to ViT and DeiT (53.58 and 51.99). This reinforces prior findings (Aggarwal et al. 2024) that the primary signal of hatefulness often resides in the caption, not the image.
These findings corroborate that counterfactual data augmentation enhances model robustness, especially in F1, by reducing overfitting to superficial features and encouraging deeper semantic reasoning. Such augmentation strategies are particularly effective when models have sufficient capacity to leverage subtle distinctions between hateful and non-hateful content, in larger contextual models like RoBERTa and CLIP.
Quality of Augmented Data
Scaled Labels’ Evaluation
To assess the reliability of the label augmentation strategy, we conduct a human evaluation on a randomly selected sample of 200 memes. Although the newly scaled labels are checked against associated hate labels and passed through a post-filtering step, we ask human annotators to independently indicate whether they agree or disagree with each assigned label. The agreement rate comes out to be 89%, indicating high consistency between human judgment and the augmented labels.
Non-Hateful Memes’ Evaluation
We also evaluate the quality of counterfactually neutral (non-hateful) memes generated through augmentation. Another human evaluation is performed on 200 randomly selected meme pairs, each consisting of an original hateful meme and its corresponding non-hateful augmented version. Annotators confirm that all original memes are hateful due to their captions, not the background images, validating our design hypothesis.
Augmented memes in the sample are evaluated on four key dimensions, rated on a 0–5 Likert scale:
  • Formatting Quality: Whether the text is placed appropriately and the meme appears complete and coherent.
  • Background Alignment: Whether the new background stays consistent with the original in tone and content, avoiding drastic shifts that may introduce bias.
  • Caption Alignment: How well the caption aligns semantically and visually with the new background.
  • Overall Quality: An overall assessment of the meme’s coherence, quality, and naturalness.
Figure 2 shows the distribution of these quality scores. In 7 out of the 200 cases, the multimodal generation agent fails to overlay the caption onto the image (no caption alignment score is assigned).
We have added a separate error analysis section with demonstrated examples in SI.
Figure 2: Quality scores of augmented non-hateful memes.
Implications
The 89% agreement rate confirms the reliability of scaled labels, supporting their use for capturing nuanced hatefulness beyond binary classification. Human evaluations of the counterfactually neutral memes validate our augmentation pipeline, showing high quality across multiple dimensions. While some errors remain, the approach offers a scalable method for generating diverse, bias-reducing training data. This supports broader efforts to build more robust and fair multimodal systems.
Conclusion
In this paper, we present a dual-pronged framework to advance multimodal hate detection through prompt optimization and multimodal data augmentation. We demonstrate that how we frame the task (via prompts) and what data we train on (via counterfactual augmentation) are critical decisions driving robust model performance.
Our prompt optimization experiments—centered on InternVL2—show that prompt structure and label granularity significantly influence outcomes. InternVL2 achieves the highest F1-scores in different settings under both fine-tuning and in zero-shot prompting. These also offer a more context-sensitive foundation for content moderation systems.
Complementing this, we introduce a novel multi-agent pipeline that generates 2,479 counterfactually neutral memes to debias models’ learning to identify hate. This is, to our knowledge, the first large-scale demonstration of counterfactual meme augmentation for hate detection. The extended dataset leads to consistent F1 improvements, particularly in models capable of deeper semantic reasoning. These gains are most pronounced in handling edge cases and class imbalance, affirming the value of data-centric strategies.
Together, our contributions break new ground by offering an integrated framework that combines prompt design with systematic data augmentation for hateful content detection. Future extensions could explore addressing the limitations of this work, such as its focus on text-centric hate, and finding computationally inexpensive alternatives for multimodal data augmentation. The generalizability of our findings is also confined to a single dataset. While we offer an ethical framework for debiasing hate datasets, the ethical implications of using generative models for content modification may require further regulatory scrutiny.
Acknowledgment
This work is supported by the Singapore Ministry of Education AcRF TIER 3 Grant (MOE-MOET32022-0001). We greatly thank Insyirah Mujtahid, Shaz Furniturewala, and Jingwei Gao for their helpful input and assistance during various stages of this work.
Appendix A Supplementary Information
Appendix B Prompt Optimization
Prompting Components
We organize all prompting strategies in this paper into modular components (see Table 4). The simple module poses direct questions to the model. The category module breaks the task into hatefulness subcategories. The scale module requires numerical outputs (e.g., 0–9), while the binary module expects true/false responses. To reduce conversion errors, we apply output constraints within the scale or binary groups. Although some other strategies perform better, we restrict prompts to clean combinations of modules to control for confounders.
Detailed Experimental Settings
The experiments are organized along three dimensions. The first dimension concerns the model category: multi-modal prompting on a large pre-trained model, InternVL2, multi-modal fine-tuning of the same model, and unimodal fine-tuning on a smaller textual model, DistilBERT. To simulate typical computational constraints, we are using LoRA, a low-rank adaptation technique, to reduce the computational overhead during fine-tuning of 8B models.
The second dimension relates to prompting strategies: either a basic prompt designed for the classification task (prompt component: simple) or a more elaborate prompt outlining specific categories of hatefulness (prompt component: category; see Table 4). The third dimension addresses output format: either binary or scaled outputs. The binary label asks for a direct classification (i.e., True or False; label component Binary), while the scale-based format captures more nuanced levels of hatefulness (label component Scale). All component settings are kept consistent across combinations for fair comparison.
As the original dataset does not include scale-based labels, we use GPT-4o-mini as a teacher model to generate them on a sample. To ensure label accuracy, we manually filter entries by cross-checking with the binary ground-truth. For example, instances labeled hateful in the ground-truth but rated as 1 (not hateful) on the 0–9 scale have been discarded. This ensures correctness while enriching the dataset with scale-level and multi-modal information from the teacher model. After this, we end up with approximately 2K instances for training purposes.
For InternVL2, we use its default fine-tuning hyperparameters. For DistilBERT, the following training configuration is applied uniformly: the learning rate is set to $2\times 10^{-5}$, with a batch size of 12 for both training and evaluation per device, trained for 12 epochs. We use a weight decay of 0.01, with evaluation and saving strategies set to occur at every epoch, and enabled loading of the best model at the end.
Appendix C Multimodal Augmentation
Prompting Components
We employ multiple agents in a structured manner for the multimodal data augmentation pipeline. The prompts and generation parameters are summarized in Table 5, with other parameters set to the models’ default values. For generating image descriptions and analyzing hatefulness in background and caption text, we use the open-source models InternLM-XComposer2.5-7B and Qwen2.5-14B, respectively. For generating new non-hateful captions and meme regeneration, we utilize models from OpenAI and Gemini, which are closed-source. These models are selected due to the complexity of the tasks and their state-of-the-art performance.
Detailed Experimental Settings
The experiments are conducted in three main directions: two unimodal (text and vision separately) and one multimodal. We evaluate the effectiveness of the extended dataset with non-hateful augmented memes. For all experiments, state-of-the-art encoder models (base variant) are fine-tuned on two types of training data. As the addition of non-hateful memes increases data imbalance, we apply downsampling, using a random seed of 42 for all processes, including training and checkpoint loading, to ensure reproducibility. Only the last two layers of each encoder are fine-tuned to prevent overfitting. Training is performed for up to 100 epochs with early stopping (patience set to 3, using weighted-F1 score on the validation set as the criterion), a batch size of 128, a learning rate of $2\times 10^{-5}$, and the AdamW optimizer to minimize cross-entropy loss.
Appendix D Error Examples in Augmented Data
To assess the effectiveness of our two data augmentation strategies: (1) scaling binary labels to continuous hatefulness scores and (2) generating new multimodal hate examples, we conducted a comprehensive human evaluation of the resulting datasets.
For the scaled label augmentation, we observe that approximately 11% of the examples yielded human evaluations that disagreed with the predicted hatefulness scores. As shown in Table 6, several of these disagreements stem from indirect or context-dependent hate speech, which can be subtle and difficult for even advanced models like GPT-4o-mini to accurately evaluate. For example, memes involving racial undertones or cultural stereotyping (e.g., Examples 1 and 3) may appear humorous on the surface but carry implicit offensive content. This highlights the challenge of detecting implicit and contextual hate, and suggests that better modeling of sociocultural nuance is needed.
In the case of multimodal data augmentation (Table 7), we identified several quality issues across a small subset of the generated samples:
  • In 7 out of 200 cases, the output image lacks any caption, although the visual background was still appropriate and non-hateful (like Example 1).
  • Example 2 retains a background similar to the original hateful meme, but the mismatch between the visual context and the newly added non-hateful caption results in a low-quality augmentation.
  • Some samples exhibit formatting errors, such as misspellings or unnatural alignment of image and text. This can also be seen in Example 2.
  • In case of Example 3, the background context is lost. The original meme features a person with a disability, while the generated version failed to preserve this detail, resulting in semantic drift.
  • Another sample shows (Example 4) a directional inconsistency in a character’s face, which disrupts the visual coherence and makes the meme appear unnatural.
These findings indicate that while automated augmentation pipelines are useful, manual or semi-automated validation remains crucial to ensure high-quality, semantically faithful, and socially responsible outputs. Furthermore, context preservation and visual-textual coherence are recurring challenges that deserve further attention.
Dataset Prompt Example
{{Simple}} + {{Binary}} {{Your task is to analyze this given image and its caption to identify if there’s any forms of hateful content.}} {{If it’s hateful, return the ‘TRUE’ — ‘FALSE’.}}
{{Simple}} + {{Scale}} {{Simple}} {{You will be provided with the image, based on this label and the content of the image, you will assign a hatefulness score on a scale from 0 to 9, where: 0  The image is not hateful at all.9  The image is extremely hateful. Return the score as an integer in range 0 to 9.}}
{{Category}} + {{Binary}} {{Simple}}{{Try to focus on the presence of any element that relates to any of the following: 1. Sexual aggression: a. Homophobia and Transphobia: This category encompasses hate speech targeting LGBTQ+ individuals, including slurs, derogatory comments, and rhetoric that seeks to undermine or dehumanize people based on their sexual orientation or gender identity. b. Misogyny and Sexism: This category includes hate speech directed at women or based on gender. It covers derogatory language, stereotypes, and rhetoric that perpetuate gender inequality, objectification, and violence against women. 2. Hate based on ideology: a. Political Hate Speech: This category includes hate speech that is politically motivated, often targeting individuals or groups based on their political beliefs or affiliations. It may include inflammatory language, threats, and rhetoric designed to polarize or incite violence within political contexts. 3. Racism and xenophobia: a. COVID-19 and Xenophobia: This category includes hate speech that arose during the COVID-19 pandemic, often targeting specific ethnic groups or nationalities. It includes xenophobic language blaming certain groups for the spread of the virus, as well as fear-mongering and scapegoating related to the pandemic. b. Racism Against Black People: This category focuses on hate speech directed at Black individuals or communities. It includes racial slurs, stereotypes, dehumanization, and other forms of derogatory language that perpetuate racial discrimination and inequality. c. Racist Hate Against Other Ethnic Groups: This category includes hate speech directed at various ethnic groups other than Black individuals. It covers a range of racial slurs, xenophobic language, dehumanization, and derogatory remarks targeting specific ethnicities or nationalities. d. White Supremacy: This category includes hate speech promoting white supremacist ideology, often intertwined with Christian extremist views. It includes rhetoric that advocates for racial superiority, anti-immigrant sentiments, and the dehumanization of non-white groups, sometimes using religious justifications for these beliefs. 4. Bigotry: a. Anti-Muslim and Islamophobic Hate: This category comprises hate speech aimed at Muslims and Islam. It includes language that promotes fear, hatred, dehumanization, or prejudice against Muslim individuals or communities, often using religious and cultural references to incite hostility. b. Anti-Semitic Hate: This category focuses on hate speech directed at Jewish people and Judaism. It includes references to historical anti-Semitic tropes, conspiracy theories, and other forms of rhetoric that seek to dehumanize or discriminate against Jewish individuals and communities. 5. Miscellaneous Hate Speech: This category captures various forms of hate speech that do not fit neatly into the other categories. It includes a wide range of derogatory language and expressions that target individuals or groups based on different aspects of identity or personal characteristics. This category includes hate speech that targets individuals based on their physical or mental disabilities. It often includes derogatory language that mocks or devalues people with disabilities, promoting harmful stereotypes and exclusion.}} {{Binary}}
{{Category}} + {{Scale}} {{Simple}}{{Category}}{{Scale}}
  • Note. a) Brackets are only used for clarification; b) each color represents a unique component of prompts.
Table 4: Prompts used for prompt optimization framework experiment.
Pipeline Component Model Generation Params* Prompt Used
Background Description
Generation
InternLM-XComposer2.5-7B
do_sample=False,
num_beams=3,
top_p=None,
use_meta=True
Analyze and describe the given image’s visuals in short. Do not include the description about the text written on the image.
Hate Component
Analyser
Qwen2.5-14B max_new_tokens=1 You are a social media content moderator. You are given a text caption of a meme and your task is to identify whether it’s hateful? Answer in ‘Yes’ or ‘No’ only. >>>Image Caption: <TEXT CAPTION> >>>isHateful:
You are a social media content moderator. You are given a text description of the background image of a meme and your task is to identify whether it’s hateful? Answer in ‘Yes’ or ‘No’ only. >>>Image Description: <IMAGE DESCRIPTION> >>>isHateful:
New Non-hateful
Caption Generation
GPT-4o-mini temperature=0 You are a helpful social media expert with deep knowledge of memes and inclusive language. Your task is to rewrite a hateful meme caption so that it becomes non-hateful and bias-reducing, while staying relevant to the background image description. The revised caption must: - Preserve reference to the originally targeted group or person (e.g., Muslims, Women, Immigrant, etc.), if they are mentioned. The new caption should try to retain the word related to target group in a respectful and bias-correcting way. - Transform the message to be neutral or positively framed toward the mentioned group/person. - Help counteract harmful stereotypes or biases. - Remain coherent and contextually appropriate for the image. Strict Rule: Only return the revised caption. Do not include explanations, disclaimers, or any other text. Meme background image description: ‘<IMAGE DESCRIPTION>’ Meme caption text: ‘<TEXT CAPTION>’ Revised non-hateful meme caption:
Meme Image
Regeneration
Gemini-Flash-2.0-exp
temperature=0,
max_output_tokens=100,
response_modalities=[“image”,“text”],
safety_settings=“BLOCK_NONE”,
response_mime_type=“text/plain”
You are a creative social media expert with deep knowledge of internet memes. Your task is to generate a humorous and engaging meme image that is non-hateful. Use the given background description and overlay the provided text in a visually appealing way. Background image description: ‘<IMAGE DESCRIPTION>’ Meme text: ‘<NEW NON-HATEFUL TEXT CAPTION>’ Ensure the meme aligns with popular internet humor while remaining appropriate and widely relatable.
Table 5: Prompts used for multimodal data augmentation pipeline. Note: Any additional parameters are set to their default values.
Example 1 Example 2 Example 3 Example 4
Meme
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
HS
6 2 6 0
Table 6: Low-quality scaled label examples (human evaluator disagrees with the hatefulness score) from the output label augmentation approach. HS refers to Hatefulness Score.
Example 1 Example 2 Example 3 Example 4
Hateful Meme
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Non-hateful Meme
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 7: Low-quality input-output examples from multimodal data augmentation pipeline.